OpenStudy (anonymous):

A boat traveled 108 miles dwnstream and back. The trip downstream took 3 hours. The trip back took 6 hours. What is the speed of the boat in still water? what is the speed of the current?

OpenStudy (anonymous):

melissa........... no.

hero (hero):

I hate these "downstream and back" questions because I don't know if they mean 108 miles total trip or 216 miles total trip.

hero (hero):

According to one unverified source, they mean 216 miles total trip so 108 miles there and 108 miles back.

OpenStudy (anonymous):

d=s1*t1 d=s2*t2 where s1=s+x s2=s-x find s

hero (hero):

Going by that... d = rt 108 = (b+w)*3 108 = (b-w)*6 108/3 = b + w 108/6 = b - w 36 = b + w 18 = b - w w = 36 - b w = b - 18 w = w 36 - b = b - 18 2b = 54 b = 54/2 b = 27 w = 9

hero (hero):

I should have used c for current instead of w for water

OpenStudy (callisto):

actually , starting from 36 = b + w -(1) 18 = b - w - (2) you can use elimination to find the unknown [ by (1) + (2)]

hero (hero):

yah. I felt like doing it my way. I hope you don't have an issue with it.

hero (hero):

I like to make my solutions look long that way it looks like it takes a lot of work to get to the solution

hero (hero):

@mimi_x3 came online

OpenStudy (mimi_x3):

lol, yess..??

hero (hero):

Oh sorry, did I say that outloud?

OpenStudy (mimi_x3):

lol, very loud that i was able to receive it that far.

hero (hero):

You must be on the other side of the world. I didn't know my voice could be that loud. But then again, God can hear me from another universe so I'm not surprised.

OpenStudy (mimi_x3):

Lol, how would you know that god exist? If he does I doubt that he would able to hear a hero.

hero (hero):

Me and God are cool. I helped him out a few times. If you want, I can introduce you to him one day. But be careful, he doesn't like people looking at his face. Kinda shy.

OpenStudy (mimi_x3):

Lol, you are hallucinating..under delirium.

OpenStudy (callisto):

just provide another way to do it only

hero (hero):

I know it's hard for you to believe.

OpenStudy (mimi_x3):

Put faith in tangible results... You can't even believe it..infact no one can.

hero (hero):

Mimi, I don't live by faith. I live by fact. My friendship with God is more valid than math.

OpenStudy (mimi_x3):

Hm, let's stop this..i don't want to get suspended.

hero (hero):

Don't worry Mimi, If you get suspended, that's more time you'll be able to spend with God

hero (hero):

God doesn't have a face, so I don't have to worry about facing anything lol

OpenStudy (mimi_x3):

lol, interpreting words in a ridiculus way..

hero (hero):

I'm about to delete all of my comments